OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x3 + 4 and g(x) = 3 square root x+4 Explain

OpenStudy (anonymous):

\[f(x)=x^3+4\] \[g(x)=\sqrt[3]{x+4}\] is there maybe a typo there?

OpenStudy (anonymous):

one should have \(-4\) i think